# Pruning via Merging: Compressing LLMs via Manifold Alignment Based Layer Merging

### v1

@@ Abstract
While large language models (LLMs) excel in many domains, their complexity and scale challenge deployment in resource-limited environments. Current compression techniques, such as parameter pruning, often fail to effectively utilize the knowledge from pruned parameters. To address these challenges, we propose Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA), a novel approach that uses manifold learning and the Normalized Pairwise Information Bottleneck (NPIB) measure to merge similar layers, reducing model size while preserving essential performance. We evaluate MKA on multiple benchmark datasets and various LLMs. Our findings show that MKA not only preserves model performance but also achieves substantial compression ratios, outperforming traditional pruning methods. Moreover, when coupled with quantization, MKA delivers even greater compression. Specifically, on the MMLU dataset using the Llama3-8B model, MKA achieves a compression ratio of 43.75% with a minimal performance decrease of only 2.82%. The proposed MKA method offers a resource-efficient and performance-preserving model compression technique for LLMs.
Pruning via Merging: Compressing LLMs via
Manifold Alignment Based Layer Merging
[…]
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI et al. (2024), Llama-3111https://github.com/meta-llama/llama3 , Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2024), have demonstrated remarkable proficiency in language understanding and generation. These models, with billions of parameters trained on trillions of tokens, can handle complex tasks and exhibit emergent abilities Brown et al. (2020); Chowdhery et al. (2023). While these models have achieved unprecedented success, their growing complexity and scale have brought to the fore significant challenges in terms of computational resources, memory requirements, and energy consumption Bender et al. (2021); Bommasani et al. (2021), raising concerns about their sustainability.
@@
To mitigate these challenges, researchers have developed various model compression techniques in LLM to reduce its parameter size while preserving performance Cheng et al. (2017); Deng et al. (2020); Ganesh et al. (2021); Zhu et al. (2023). These techniques can be roughly categorized into two main mainstreams Men et al. (2024): quantization Gholami et al. (2021); Li et al. (2024); Dettmers et al. (2022); Gong et al. (2024); Li et al. (2024) and pruning LeCun et al. (1989); Han et al. (2016); Gupta and Agrawal (2022); Ma et al. (2023a). Quantization based methods aid in the reduction of the memory consumption of weights, activations, and KV caches by using the low-precision values with fewer bits instead of the high-precision values. However, the acceleration benefits of quantization are seriously dependent on hardware support Tao et al. (2023) and sometimes require additional fine-tuning to maintain performance Dettmers et al. (2023); Men et al. (2024). Compared to quantization, pruning, especially structural pruning Li et al. (2017), eliminates redundant LLM’s parameters to decrease the overall parameter count, and can be applied directly to a trained LLM without retraining and is generally more hardware-friendly than quantization approaches. While effective, pruning usually risks losing valuable model structures and determining how to prune the LLM with minimal disruption to the origin remains an unsolved problem Ma et al. (2023b).
@@
Figure 1: Manifold-Based Knowledge Alignment and Layer Merging (MKA) framework consists of two main components: (1) The left side illustrates manifold learning for LLM knowledge extraction, where layer activations are transformed into low-dimensional manifolds using the Diffusion Kernel algorithm. (2) The right side depicts the similarity-based layer merging process, employing the NPIB metric to identify layers with aligned knowledge.
@@
To tackle this issue head-on, we delve into the realm of model merging Wortsman et al. (2022), a powerful technique that seamlessly weaves together the strengths and knowledge of multiple models, creating a robust and efficient aggregation. This technique, through averaging the weights of multiple models with the same architecture, can retain essential features without significant additional resources Liu et al. (2024); Wan et al. (2024). Furthermore, by offsetting the biases and errors of individual models, model merging often leads to greatly improved performance Li et al. (2023). Additional, the number of models in the merging process can be gradually and naturally reduced. However, such a useful technology are limited to merging between models currently, and few studies pay attention on merging the same internal structures within a model.
@@
Manifold Learning for LLM Knowledge: We employ manifold learning techniques to align knowledge across layers by extracting layer activations from a LLM and applying the Diffusion Kernel algorithm Tenenbaum et al. (2000) to learn low-dimensional manifold representations. This approach captures the nonlinear structure in the activation and achieves dimensionality reduction while preserving important activation features, enabling more effective comparison of knowledge patterns across different layers.
Similarity Alignment Layer Merging: Following manifold learning, we use the Normalized Pairwise Information Bottleneck (NPIB) measure Tishby et al. (2000) to construct a similarity matrix that quantifies the similarity between layers by maximizing their mutual information while considering the entropy of each layer. Based on this similarity matrix, we select the most similar layer pairs for merging.
@@
  • We develop a manifold-based knowledge alignment approach, utilizing the Diffusion Kernel and Normalized Pairwise Information Bottleneck (NPIB) to effectively capture and align similarities between layers in the parameter space.
@@
2 Manifold-Based Knowledge Alignment and Layer Merging
Our MKA method relies on the redundancy present in the latter layers of post-training LLMs Gromov et al. (2024). By merging layers with high input-output similarity from back to front, we maintain the model’s performance while reducing its size. In this section, we first describe the extraction and dimensionality reduction processes for the intermediate states, as high-dimensional intermediate states are challenging to analyze. Then, we propose our layer merging method based on similarity alignment, which aims to maintain performance by aligning intermediate states through merging techniques.
2.1 Manifold Learning for LLM Knowledge
To effectively align knowledge across LLM’s layers, MKA employs manifold learning techniques that can capture the intricate nonlinear dependencies within the LLM’s internal structure. This approach allows us to compare and align layer activations in a meaningful way, preserving essential information while reducing model complexity.
The process begins with the extraction of layer activations ${H}^{l}$ from a LLM on the dataset $\omega$ These activations represent the outputs of each layer given a set of input samples, encapsulating the knowledge learned at different stages. To transform these high-dimensional activations into a lower-dimensional space that preserves their essential features and geometric structure, we apply the Diffusion Kernel algorithm Coifman and Lafon (2006). Here are the key steps involved in this process:
Extracting Layer Activations: For each layer $l$, we extract the activations ${H}^{l}$ given input samples. These activations $\mathbf{H}^{l}$ are computed using the following equation:
$\mathbf{H}^{l}$ $=\text{LayerNorm}\left(\mathbf{H}^{l-1}+\text{MultiHead}\left(\mathbf{H}^{l-1}\right)\right)$
$\quad+\text{FeedForward}\left(\mathbf{H}^{l-1}\right)$ (1)
Constructing the Pairwise Distance Matrix: Next, we calculate the pairwise Euclidean distance matrix $D$ for the activations ${H}^{l}$. This matrix captures the distances between all pairs of activations, serving as the basis for the manifold learning process.
Applying the Diffusion Kernel: We apply the Diffusion Kernel to transform the distance matrix $D$ into low-dimensional manifold representations $\Phi_{i}$, capturing the intrinsic geometric structure of the data. The kernel function smooths the data, emphasizing the intrinsic geometric structure:
$\mathbf{E}$ $=\text{EigVectors}_{d}\left(\text{Diag}\left(\sum_{j}e^{-\left(\frac{\|\mathbf{H}_{i}-\mathbf{H}_{j}\|^{2}}{\sigma K}\right)^{0.5}}\right)\right.$
$\quad\left.-e^{-\left(\frac{\|\mathbf{H}_{i}-\mathbf{H}_{j}\|^{2}}{\sigma K}\right)^{0.5}}\right)$ (2)
where $\sigma_{K}$ is the kernel bandwidth parameter, and $\text{EigVectors}_{d}$ refers to the eigenvectors corresponding to the $d$ smallest eigenvalues of the Laplacian matrix $L$. This transformation captures the essential features and relationships within the activations, enabling effective comparisons across different layers.
2.2 Similarity-based Layer Merging

### v2

@@ Abstract
While large language models (LLMs) excel in many domains, their complexity and scale challenge deployment in resource-limited environments. Current compression techniques, such as parameter pruning, often fail to effectively utilize the knowledge from pruned parameters. To address these challenges, we propose Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA), a novel approach that uses manifold learning and the Information Bottleneck (IB) measure to merge similar layers, reducing model size while preserving essential performance. We evaluate MKA on multiple benchmark datasets and various LLMs. Our findings show that MKA not only preserves model performance but also achieves substantial compression ratios, outperforming traditional pruning methods. Moreover, when coupled with quantization, MKA delivers even greater compression. Specifically, on the MMLU dataset using the Llama3-8B model, MKA achieves a compression ratio of 43.75% with a minimal performance decrease of only 2.82%. The proposed MKA method offers a resource-efficient and performance-preserving model compression technique for LLMs. We make our code available at https://github.com/SempraETY/Pruning-via-Merging
Index Terms:
Model Compression, Layer Merging, Manifold Learning, Large Language Models (LLMs)
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 [1], Llama-3 [2], Llama-2 [3] and Mistral [4], have demonstrated remarkable proficiency in language understanding and generation. These models, with billions of parameters trained on trillions of tokens, can handle complex tasks and exhibit emergent abilities [5, 6]. While these models have achieved unprecedented success, their growing complexity and scale have brought to the fore significant challenges in terms of computational resources, memory requirements, and energy consumption [7, 8], raising concerns about their sustainability.
@@
To mitigate these challenges, researchers have developed various model compression techniques in LLM to reduce its parameter size while preserving performance [9, 10, 11, 12, 13]. These techniques can be roughly categorized into two main mainstreams [14]: quantization [15, 16, 17, 18, 16] and pruning [19, 20, 21, 22]. Quantization based methods aid in the reduction of the memory consumption of weights, activations, and KV caches by using the low-precision values with fewer bits instead of the high-precision values. However, the acceleration benefits of quantization are seriously dependent on hardware support [23] and sometimes require additional fine-tuning to maintain performance [24, 14]. Compared to quantization, pruning, especially structural pruning [25], eliminates redundant LLM’s parameters to decrease the overall parameter count, and can be applied directly to a trained LLM without retraining and is generally more hardware-friendly than quantization approaches. While effective, pruning usually risks losing valuable model structures and determining how to prune the LLM with minimal disruption to the origin remains an unsolved problem [26].
@@
Figure 1: Manifold-Based Knowledge Alignment and Layer Merging (MKA) framework consists of two main components: (1) The left side illustrates manifold learning for LLM knowledge extraction, where layer activations are transformed into low-dimensional manifolds using the Diffusion Kernel algorithm. (2) The right side depicts the similarity-based layer merging process, employing the IB metric to identify layers with aligned knowledge.
@@
To tackle this issue head-on, we delve into the realm of model merging [27], a powerful technique that seamlessly weaves together the strengths and knowledge of multiple models, creating a robust and efficient aggregation. This technique, through averaging the weights of multiple models with the same architecture, can retain essential features without significant additional resources [28, 29]. Furthermore, by offsetting the biases and errors of individual models, model merging often leads to greatly improved performance [30]. Additional, the number of models in the merging process can be gradually and naturally reduced. However, such a useful technology are limited to merging between models currently, and few studies pay attention on merging the same internal structures within a model.
@@
Manifold Learning for LLM Knowledge: We employ manifold learning techniques to align knowledge across layers by extracting layer activations from a LLM and applying the Diffusion Kernel algorithm [31] to learn low-dimensional manifold representations. This approach captures the nonlinear structure in the activation and achieves dimensionality reduction while preserving important activation features, enabling more effective comparison of knowledge patterns across different layers.
Similarity Alignment Layer Merging: Following manifold learning, we use the Information Bottleneck (IB) measure [32] to construct a similarity matrix that quantifies the similarity between layers by maximizing their mutual information while considering the entropy of each layer. Based on this similarity matrix, we select the most similar layer pairs for merging.
@@
  • We develop a manifold-based knowledge alignment approach, utilizing the Diffusion Kernel and Information Bottleneck (IB) to effectively capture and align similarities between layers in the parameter space.
@@
2 Related Work
Our proposed Manifold-Based Knowledge Alignment and Layer Merging (MKA) framework builds upon and integrates several key areas of research in deep model optimization for efficiency. In the following subsections, we first explore methods for learning high-dimensional data, which are essential for understanding the complex structures within large language models. We then review the primary model compression techniques, including quantization, pruning, and knowledge distillation, which aim to reduce the computational and memory footprint of these models. Finally, we delve into the emerging field of model merging, highlighting recent advancements that inform our approach to aggregating and aligning knowledge across model layers.
2.1 Methods for Learning High-Dimensional Data.
High-dimensional data poses significant challenges in machine learning, primarily due to the curse of dimensionality, which can lead to increased computational complexity and overfitting [33]. To address these challenges, manifold learning techniques have been extensively developed to capture the underlying low-dimensional structures within high-dimensional datasets. Techniques such as Principal Component Analysis (PCA) [34], t-Distributed Stochastic Neighbor Embedding (t-SNE) [35], and Uniform Manifold Approximation and Projection (UMAP) [36] have been widely adopted for dimensionality reduction and data visualization. More recently, diffusion-based methods like Diffusion Maps [37] and the Diffusion Kernel algorithm [31] have shown promise in preserving the intrinsic geometry of data while facilitating efficient computation. These manifold learning approaches are crucial for understanding and aligning complex patterns in high-dimensional spaces, making them well-suited for applications in large language model (LLM) compression and knowledge alignment.
2.2 Model Compression Methods.
Model compression has emerged as a vital area of research aimed at reducing the computational and memory footprint of large-scale models without significantly compromising their performance [9, 10, 11, 12, 13]. The primary techniques in model compression include quantization, pruning, and knowledge distillation.
Quantization involves reducing the precision of the model’s weights and activations from high-bit representations (e.g., 32-bit floating-point) to lower-bit formats (e.g., 8-bit integers), thereby decreasing memory usage and accelerating inference [15, 16, 17, 18]. While effective, quantization often requires specialized hardware support and may necessitate additional fine-tuning to maintain model accuracy [24, 14].
Pruning techniques focus on eliminating redundant or less significant parameters from the model, thereby reducing the overall parameter count and computational requirements [19, 20, 21, 22]. Pruning can be categorized into unstructured pruning, which removes individual weights, and structured pruning, which removes entire neurons or channels [25]. Structural pruning is particularly advantageous for its hardware friendliness and the ability to directly apply it to pre-trained models without extensive retraining [26].
Knowledge Distillation involves training a smaller "student" model to replicate the behavior of a larger "teacher" model, thereby transferring knowledge while reducing model size [38]. This approach has been effective in maintaining performance levels while achieving significant compression.
Recent advancements have also explored hybrid approaches that combine multiple compression techniques to leverage their complementary strengths [14]. These methods aim to achieve higher compression ratios and better performance retention by integrating quantization, pruning, and distillation strategies.
2.3 Model Merging.
Model merging is a burgeoning area that focuses on combining multiple models to create a single, more robust model by leveraging the strengths and knowledge of its constituents [27]. Traditional approaches, such as Model Soup [27], utilize simple weight averaging to merge models with the same architecture, effectively blending their learned representations. However, this method can be limited by the need for identical architectures and the potential for performance degradation if the individual models are not sufficiently aligned.
Advancements in model merging have introduced more sophisticated techniques to enhance the robustness and performance of the merged models. Checkpoint Merging [28] employs Bayesian optimization to selectively weight and integrate different model checkpoints, resulting in a more stable and high-performing merged model. Similarly, MindMerger [39] facilitates the fusion of models with varying specializations, thereby enhancing the overall capabilities of the resultant model by integrating diverse knowledge bases.
Dynamic expert merging methods, such as DELLA-Merging [40], dynamically incorporate specialized expert models during inference, allowing the merged model to adapt to a wide range of tasks. Adaptive weighting approaches like AdaMerging [41] and MetaGPT [42] leverage meta-learning and dynamic weighting schemes to fine-tune the merging process, ensuring optimal integration of the strengths of constituent models.
Furthermore, task-oriented merging strategies, including Task Arithmetic [43], Language and Task Arithmetic [44], and Task Arithmetic in Tangent Space [45], focus on blending models trained on different tasks to create versatile LLMs capable of handling multiple applications. These approaches enhance the flexibility and applicability of merged models, making them more adaptable to diverse real-world scenarios.
Despite the progress in model merging, most existing methods focus on merging distinct models rather than addressing the internal layer structures within a single model. Our proposed Manifold-Based Knowledge Alignment and Layer Merging (MKA) framework addresses this gap by enabling the compression of LLMs through the progressive aggregation of knowledge across layers, thereby reducing the total number of layers while preserving essential model performance.
3 Preliminaries
In this section, we introduce the foundational concepts and theoretical frameworks essential for understanding the proposed Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA) method. We begin by discussing overparameterization and redundancy in Large Language Models (LLMs), then present the manifold hypothesis in the context of neural representations and diffusion geometry, and finally introduce mutual information within the Information Bottleneck (IB) framework.
3.1 Overparameterization and Redundancy in Large Language Models
Let $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ denote a dataset, where $\mathbf{x}_{i}\in\mathcal{X}$ are input sequences and $y_{i}\in\mathcal{Y}$ are the corresponding targets. A Large Language Model (LLM) is modeled as a parameterized function $f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}$ with parameters $\bm{\theta}\in\mathbb{R}^{d}$, where $d$ is the total number of parameters. An LLM is said to be overparameterized if $d\gg N$. Overparameterization often results in redundancy; that is, different parameter configurations yield similar functional mappings. Formally, for a given $\epsilon>0$ and small constant $\delta$, a subset of parameters $\mathcal{S}\subset\bm{\theta}$ is considered redundant if there exists an alternative parameter set $\bm{\theta}^{\prime}$ satisfying
$\|\bm{\theta}^{\prime}-\bm{\theta}\|\leq\epsilon\quad\text{and}\quad L(f_{\bm{\theta}^{\prime}},\mathcal{D})\leq L(f_{\bm{\theta}},\mathcal{D})+\delta.$ (1)
This observation motivates the exploration of layer merging as a strategy for model compression.
3.2 Manifold Hypothesis in Neural Representations and Diffusion Geometry
The manifold hypothesis posits that high-dimensional data encountered in real-world applications lie on a low-dimensional manifold embedded in the ambient space $\mathbb{R}^{D}$. A manifold $\mathcal{M}$ of dimension $k$ is a topological space such that for every point $p\in\mathcal{M}$, there exists a neighborhood $U$ and a homeomorphism $\phi:U\rightarrow\mathbb{R}^{k}$.
In the context of neural networks, let $\mathbf{H}^{l}_{i}\in\mathbb{R}^{d_{l}}$ denote the activation of layer $l$ for input $\mathbf{x}_{i}$. The collection $\{\mathbf{H}^{l}_{i}\}_{i=1}^{N}$ is assumed to lie on a manifold $\mathcal{M}_{l}$ of intrinsic dimension $k_{l}$ (with $k_{l}\ll d_{l}$) embedded in $\mathbb{R}^{d_{l}}$. This assumption is supported by the observation that deep networks implicitly perform dimensionality reduction.
To analyze the intrinsic geometry of the activation manifold, we utilize diffusion geometry. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node corresponds to an activation vector and the edge weights are defined via a Gaussian kernel. Specifically, for two activation vectors $\mathbf{H}^{l}_{i}$ and $\mathbf{H}^{l}_{j}$, the affinity is given by
$W_{ij}=\exp\left(-\frac{\|\mathbf{H}^{l}_{i}-\mathbf{H}^{l}_{j}\|^{2}}{\sigma^{2}}\right),$ (2)
where $\sigma>0$ controls the neighborhood scale. The degree matrix $D$ is defined by
$D_{ii}=\sum_{j=1}^{N}W_{ij},$ (3)
and the diffusion operator is then
$P=D^{-1}W.$ (4)
The operator $P$ captures the connectivity of the data manifold by serving as the transition probability matrix for a random walk on $\mathcal{G}$.
The diffusion map is constructed via the spectral decomposition of $P$. Suppose that $P$ is reversible (which follows from the symmetry of $W$) and let $\{(\lambda_{j},\phi_{j})\}_{j=1}^{N}$ be its eigenpairs. The diffusion map at time $t$ is defined as
$\Phi_{t}(i)=\left(\lambda_{1}^{t}\phi_{1}(i),\lambda_{2}^{t}\phi_{2}(i),\dots,\lambda_{k}^{t}\phi_{k}(i)\right).$ (5)
In many cases, the first eigenvector corresponding to $\lambda_{1}=1$ is omitted to focus on the nontrivial geometry; the indexing may be adjusted accordingly.
3.3 Mutual Information and the Information Bottleneck Principle
Mutual information is used to quantify the similarity between representations. For a continuous random variable $\mathbf{X}$ with density $p(\mathbf{x})$, the differential entropy is defined as
[…]
The mutual information between $\mathbf{X}$ and $\mathbf{Y}$ is given by
$I(\mathbf{X};\mathbf{Y})=H(\mathbf{X})+H(\mathbf{Y})-H(\mathbf{X},\mathbf{Y}).$ (7)
Within our framework, mutual information serves as a metric for comparing the diffusion map embeddings of different layers.
The Information Bottleneck (IB) principle provides a framework for obtaining a compressed representation that retains relevant information. For random variables $\mathbf{X}$ and $\mathbf{Y}$, the IB objective is to find a mapping $p(\mathbf{T}|\mathbf{X})$ that minimizes
$\min_{p(\mathbf{T}|\mathbf{X})}\;I(\mathbf{X};\mathbf{T})-\beta I(\mathbf{T};\mathbf{Y}),$ (8)
where $\beta>0$ balances compression against the preservation of relevant information. In our context, the goal is to merge the diffusion map embeddings from two layers into a single representation that maximizes mutual information with the target variable $\mathbf{Y}$ while minimizing redundancy.
4 Manifold Learning for Internal Representations
In this section, we introduce the Manifold-Based Knowledge Alignment (MKA) framework, which leverages the manifold hypothesis and diffusion geometry to identify and merge redundant layers within large language models (LLMs). Building on the concepts outlined in Section 3, we describe the extraction of high-dimensional activations, the construction of the diffusion operator, the spectral decomposition leading to diffusion maps, and the subsequent alignment and merging of layers using information-theoretic measures.
4.1 Extraction of High-Dimensional Activations
Let $\mathcal{M}$ denote an LLM processing a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$. For each input $\mathbf{x}_{i}$, the model produces activations at each layer. Specifically, we denote by $\mathbf{H}^{l}_{i}\in\mathbb{R}^{d_{l}}$ the activation of layer $l$ for input $\mathbf{x}_{i}$, where $d_{l}$ is the activation dimension at layer $l$. The activations are computed via
$\mathbf{H}^{0}_{i}$ $=\mathrm{Embed}(\mathbf{x}_{i}),$ (9)
$\mathbf{H}^{l}_{i}$ $=f_{\bm{\theta}_{l}}(\mathbf{H}^{l-1}_{i}),\quad l=1,2,\dots,L,$ (10)
where $f_{\bm{\theta}_{l}}$ denotes the transformation associated with layer $l$, parameterized by $\bm{\theta}_{l}$. The collection $\{\mathbf{H}^{l}_{i}\}_{i=1}^{N}$ thus forms a high-dimensional dataset intrinsic to each layer.
4.2 Construction of the Diffusion Operator
To analyze the manifold structure underlying these activations, we construct a weighted graph $\mathcal{G}_{l}=(\mathcal{V}_{l},\mathcal{E}_{l})$ for each layer $l$, where each node corresponds to an activation vector $\mathbf{H}^{l}_{i}$. The edges are weighted according to the affinity between activation vectors.
Given two activation vectors $\mathbf{H}^{l}_{i}$ and $\mathbf{H}^{l}_{j}$, we define the affinity using a Gaussian kernel:
$W_{ij}=K(\mathbf{H}^{l}_{i},\mathbf{H}^{l}_{j})=\exp\left(-\frac{\|\mathbf{H}^{l}_{i}-\mathbf{H}^{l}_{j}\|^{2}}{\sigma^{2}}\right),$ (11)
where $\sigma>0$ is a bandwidth parameter that controls the local neighborhood scale.
The degree matrix $D$ is diagonal with entries
$D_{ii}=\sum_{j=1}^{N}W_{ij}.$ (12)
We then define the diffusion operator as the normalized affinity matrix:
$P=D^{-1}W.$ (13)
This operator governs the transition probabilities of a random walk on the graph $\mathcal{G}_{l}$ and thus captures the intrinsic geometry of the activation manifold.
4.3 Spectral Decomposition and Diffusion Maps
The operator $P$ encodes the manifold geometry, and its spectral properties allow us to construct diffusion maps—lower-dimensional embeddings that preserve the manifold’s structure.
Assume that the affinity matrix $W$ is symmetric and positive semidefinite and that $P$ is reversible (which is the case when the underlying Markov chain is reversible). Although $P$ is not symmetric, it is similar to the symmetric matrix
$\widetilde{P}=D^{-1/2}WD^{-1/2}.$ (14)
Since $\widetilde{P}$ is symmetric, it has a complete set of real eigenvalues $\{\lambda_{k}\}_{k=1}^{N}$ and corresponding orthonormal eigenvectors $\{\bm{\phi}_{k}\}_{k=1}^{N}$, so that
$P\bm{\phi}_{k}=\lambda_{k}\bm{\phi}_{k}.$ (15)
Without loss of generality, we order the eigenvalues as
$1\displaystyle 1=\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}\geq-$ 1 . (16)
The largest eigenvalue is $\lambda_{1}=1$, and its corresponding eigenvector is constant.
Using the spectral decomposition, the diffusion map at time $t$ is defined as the mapping
$\Phi_{t}(\mathbf{H}^{l}_{i})=\Bigl{(}\lambda_{2}^{t}\,\phi_{2}(i),\,\lambda_{3}^{t}\,\phi_{3}(i),\,\dots,\,\lambda_{k+1}^{t}\,\phi_{k+1}(i)\Bigr{)},$ (17)
where $k$ is the target embedding dimension and $\phi_{j}(i)$ denotes the $i$th component of $\bm{\phi}_{j}$. (Note that the trivial eigenvector associated with $\lambda_{1}$ is omitted.)
4.4 Preservation of Manifold Structure
The diffusion map is designed to preserve the intrinsic geometry of the activation manifold by capturing multi-scale connectivity. In particular, the Euclidean distance in the diffusion space approximates the diffusion distance between points.
For any two activation vectors $\mathbf{H}^{l}_{i}$ and $\mathbf{H}^{l}_{j}$, the diffusion distance is defined by
$D_{t}(i,j)^{2}=\sum_{k=2}^{k+1}\lambda_{k}^{2t}\,\bigl{(}\phi_{k}(i)-\phi_{k}(j)\bigr{)}^{2}.$ (18)
By the very definition of the diffusion map,
$\|\Phi_{t}(\mathbf{H}^{l}_{i})-\Phi_{t}(\mathbf{H}^{l}_{j})\|_{2}^{2}=\sum_{k=2}^{k+1}\lambda_{k}^{2t}\,\bigl{(}\phi_{k}(i)-\phi_{k}(j)\bigr{)}^{2},$ (19)
so that the Euclidean distance in the embedding space approximates the diffusion distance. Moreover, as $t\to\infty$, the influence of the smaller eigenvalues diminishes, and the diffusion map increasingly emphasizes the global structure of the manifold.
4.5 Layer Similarity Measure via Mutual Information
To quantify the similarity between representations in different layers, we compare their diffusion map embeddings. Let
$\mathbf{\Psi}^{l}=\Phi_{t}(\mathbf{H}^{l})\quad\text{and}\quad\mathbf{\Psi}^{m}=\Phi_{t}(\mathbf{H}^{m})$ (20)
be the diffusion map embeddings for layers $l$ and $m$, respectively. We define their similarity in terms of the mutual information (MI)
$I(\mathbf{\Psi}^{l};\mathbf{\Psi}^{m})=H(\mathbf{\Psi}^{l})+H(\mathbf{\Psi}^{m})-H(\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}),$ (21)
where $H(\cdot)$ denotes differential entropy. In order to simplify the calculation, we assume that $\mathbf{\Psi}^{l}$ and $\mathbf{\Psi}^{m}$ are jointly Gaussian random variables. Under this joint Gaussianity assumption, the MI can be written in closed form as
$I(\mathbf{\Psi}^{l};\mathbf{\Psi}^{m})=\frac{1}{2}\ln\!\left(\frac{|\Sigma_{\mathbf{\Psi}^{l}}||\Sigma_{\mathbf{\Psi}^{m}}|}{|\Sigma_{\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}}|}\right),$ (22)
where $\Sigma_{\mathbf{\Psi}^{l}}$ and $\Sigma_{\mathbf{\Psi}^{m}}$ denote the covariance matrices of $\mathbf{\Psi}^{l}$ and $\mathbf{\Psi}^{m}$, respectively, and $\Sigma_{\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}}$ is their joint covariance matrix.
To facilitate comparison across different layer pairs, we further define the normalized mutual information (NMI) as
$S_{lm}=\frac{I(\mathbf{\Psi}^{l};\mathbf{\Psi}^{m})}{\sqrt{H(\mathbf{\Psi}^{l})H(\mathbf{\Psi}^{m})}}.$ (23)
This normalization constrains $S_{lm}$ to lie in a consistent range, thereby enabling meaningful similarity assessments between layers.
4.6 Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA)
Algorithm 1 summarizes the MKA procedure. In brief, for each layer the activations are first extracted and embedded using the diffusion map. Next, pairwise similarities between layers are computed via their mutual information. Finally, when the similarity score between a pair of layers exceeds a predetermined threshold $\tau$, the layers are merged according to a weighted combination of their parameters.